# Seal: Symbolic Execution with Separation Logic

### v1

@@ 2 Strengths and Weaknesses
Table 1 provides a view on the results of MemSafety-LinkedLists [4], the only base category in which Seal participates, illustrating the strengths of Seal. Since almost half of the programs in this category manipulate lists of bounded size only and can therefore be verified using bounded model checking (or similar heuristics), we focus on the subset of benchmarks that require reasoning about unbounded data structures. Specifically, we exclude the benchmark sets list-simple and list-ext3-properties, which operate exclusively on bounded lists, as well as ddv-machzwd where the list-manipulating code is unreachable. Only the tools that can verify correctness of at least one of the remaining programs are shown in the table [^1] Since PredatorHP spuriously returns UNKNOWN for all tasks in SV-COMP’26, we used its results from SV-COMP’25. The benchmarks used for comparison were the same in both years..
@@
As Table 1 shows, out of 23 verifiers participating in MemSafety, only 4 of them were able to verify correctness of programs with unbounded data structures. This indicates that reasoning about such programs remains a challenging task. Seal can verify more programs than 2ls [15, 14], but less than PredatorHP [13, 18] and CPAchecker [2, 1], two analysers based on symbolic memory graphs (SMG). We believe that the biggest strength of Seal lies in its potential to be extended for data structures that cannot be represented by SMGs, such as skip-lists or trees. Indeed, Astral in its new versions supports user-defined inductive predicates, including various kinds of trees, which may be implemented in Seal or potentially learned automatically. We would also like to use the generality of our technique to create a witness validator for correctness memory safety witnesses, once a witness format is established.
@@
All programs (134) Unbounded (69)
Analyser True False No res. Time [s] True False No res. Time [s]
PredatorHP 96 28 10 109 46 14 9 75
CPAchecker 87 27 20 2158 40 13 16 1550
SEAL 55 4 75 571 14 2 53 166
2LS 46 14 73 1142 3 13 53 96

### v2

@@ 2 Strengths and Weaknesses
All programs (134) Unbounded (69)
Verifier True False No res. Time [s] True False No res. Time [s]
PredatorHP 96 28 10 112 46 14 9 77
CPAchecker 87 27 20 2158 40 13 16 1550
Seal 55 4 75 571 14 2 53 166
2ls 46 14 73 1154 3 13 53 92
Table 1 provides a view on the results of MemSafety-LinkedLists [4], the only base category in which Seal participates, illustrating the strengths of Seal. Since almost half of the programs in this category manipulate lists of bounded size only and can therefore be verified using bounded model checking (or similar heuristics), we focus on the subset of benchmarks that require reasoning about unbounded data structures. Specifically, we exclude the benchmark sets list-simple and list-ext3-properties, which operate exclusively on bounded lists, as well as ddv-machzwd where the list-manipulating code is unreachable. Only the tools that can verify correctness of at least one of the remaining programs are shown in the table.
@@
As Table 1 shows, out of 20 tools (excluding metaverifiers) participating in MemSafety, only 4 of them were able to verify correctness of programs with unbounded data structures. This indicates that reasoning about such programs remains a challenging task. Seal can verify more programs than 2ls [14, 15], but less than PredatorHP [13, 18] and CPAchecker [2, 1], two analysers based on symbolic memory graphs (SMG). We believe that the biggest strength of Seal lies in its potential to be extended for data structures that cannot be represented by SMGs, such as skip-lists or trees. Indeed, Astral in its new versions supports user-defined inductive predicates, including various kinds of trees, which may be implemented in Seal or potentially learned automatically. We would also like to use the generality of our technique to create a witness validator for correctness memory safety witnesses, once a witness format is established.